OpenStudy (theghostlycreeper):

What is the value of this expression? (−2)5

OpenStudy (theghostlycreeper):

the 5 is smaller

OpenStudy (retireed):

-2^5 It is -2 to the 5th power it equals (-2)(-2)(-2)(-2)(-2) = ???

OpenStudy (theghostlycreeper):

what

OpenStudy (retireed):

What is (-2)(-2)(-2)(-2)(-2) = __________

OpenStudy (theghostlycreeper):

+ - x division

OpenStudy (retireed):

TIMES What is (-2) times (-2) =

OpenStudy (theghostlycreeper):

4

OpenStudy (retireed):

What is 4 times -2 = _____

OpenStudy (theghostlycreeper):

-8

OpenStudy (retireed):

What is -8 times -2 = _____

OpenStudy (theghostlycreeper):

16

OpenStudy (retireed):

What is 16 times -2 = _____

OpenStudy (theghostlycreeper):

32

OpenStudy (retireed):

Try again.

OpenStudy (theghostlycreeper):

-32

OpenStudy (retireed):

What is (-2)(-2)(-2)(-2)(-2) = __________

OpenStudy (theghostlycreeper):

-32

OpenStudy (retireed):



OpenStudy (retireed):

Also typed as -2^5 = _______

OpenStudy (theghostlycreeper):

did we not just do all that math not for the answear

OpenStudy (theghostlycreeper):

can you just tell me the answear i have 15 mins to coomplet the quiz

OpenStudy (retireed):

I don't understand the last question. You answered the question yourself.

OpenStudy (theghostlycreeper):

your a wizard

OpenStudy (retireed):

Thanks, I guess.

OpenStudy (mathmale):

Let's start over. (-2)5 has no meaning other than -2*5 = -10. I think your math problem is (-2)^5, or \[(-2)^5\]

OpenStudy (mathmale):

You MUST raise the exponent 5 above the level of the base (-2). \[(-2)^5=(-1)^5(2^5)\]

OpenStudy (mathmale):

Can you now evaluate (-1)^5? Can you evaluate 2^5?

OpenStudy (retireed):

If you read the poster's second entry ..... "the 5 is smaller" I knew he meant it was an exponent